Suppose you bought a bond with an annual coupon of 7 percent one year ago for $1,010. The bond sells for $985 today.

a.

Assuming a $1,000 face value, what was your total dollar return on this investment over the past year?

b.

What was your total nominal rate of return on this investment over the past year? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

c.

If the inflation rate last year was 3 percent, what was your total real rate of return on this investment? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a) Dollar return = Selling Price - Buying Price + Coupon

Dollar return = $985 - $1010 + $70 = $45

b) Rate of return = Dollar return/Buy price = 45/1010 = 4.45%

c) Based on Fischer relation,

(1 + Nominal rate) = (1 + Real rate) * (1 + Inflation)

(1 + 4.45%) = (1 + Real rate) * (1 + 3%)

Real rate = 1.41%